# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ 3.3. Augmented geodesic synthetic control method
Let $(\mathcal{M},d)$ be a unique geodesic space. For any two points $\alpha,\beta\in\mathcal{M}$, there exists a geodesic transport map $\Gamma_{\alpha,\beta}:\mathcal{M}\to\mathcal{M}$ satisfying $\Gamma_{\alpha,\beta}(\alpha)=\beta$ and for any $\omega\in\mathcal{M}$, there exists a unique point $\zeta\in\mathcal{M}$ such that $\Gamma_{\alpha,\beta}(\omega)=\zeta$.
This assumption ensures that the geodesic joining points $\alpha$ and $\beta$ can be naturally extended from any other point $\omega\in\mathcal{M}$, yielding a uniquely defined endpoint $\zeta\in\mathcal{M}$. In Euclidean space $\mathbb{R}^{q}$, the transport map is explicitly defined as $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$, and this definition extends naturally to general vector spaces, where geodesics correspond to vector translations, as well as to Riemannian manifolds via parallel transport (Yuan et al.,, 2012; Lin and Yao,, 2019; Chen et al., 2023a, ). Such geodesic transport maps exist widely in metric spaces encountered in practical applications (Zhu and Müller,, 2023); see Remark 3.5 for further details.
@@
Consider a metric space $(\mathcal{M}_{X},d_{X})$ that may differ from $(\mathcal{M},d)$ and the corresponding data pairs $\{(\bm{Y}_{j,1:T},\bm{X}_{j,1:T_{0}})\}_{1\leq j\leq J+1}$, where $\bm{Y}_{j,1:s}=(Y_{j,1},\dots,Y_{j,s})\in\mathcal{M}^{s}$ and $\bm{X}_{j,1:s}=(X_{j,1},\dots,X_{j,s})\in\mathcal{M}_{X}^{s}$. Denote by $m_{t}(\bm{Z}_{j,1:T_{0}})$ a regression model aiming to predict $Y_{j,t}(N)$ for $t=T_{0}+1,\dots,T$, where $\bm{Z}_{j,1:T_{0}}=(\bm{Y}_{j,1:T_{0}},\bm{X}_{j,1:T_{0}})$ and let $\widehat{m}_{t}(\cdot)$ be its estimated version using data $\{(Y_{j,t},\bm{Z}_{j,1:T_{0}})\}_{2\leq j\leq J+1}$. Define
$m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,m_{t}(\bm{Z}_{j,1:T_{0}})),$
@@
$Y_{1,t}^{(\mathrm{AGSC})}$ $=\Gamma_{\widehat{m}_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),\widehat{m}_{t}(\bm{Z}_{1,1:T_{0}})}(Y_{1,t}^{(\mathrm{GSC})}).$
To see the rationale for this estimator, denote by $\mathcal{G}(\mathcal{M})=\{\gamma_{\alpha,\beta}:\alpha,\beta\in\mathcal{M}\}$ the space of geodesics of the unique geodesic space $\mathcal{M}$. To compare geodesics, we define a binary relation $\sim$ on $\mathcal{G}(\mathcal{M})$ through the geodesic transport map, where $\gamma_{\alpha_{1},\beta_{1}}\sim\gamma_{\alpha_{2},\beta_{2}}$ if and only if $\Gamma_{\alpha_{1},\beta_{1}}(\omega)=\Gamma_{\alpha_{2},\beta_{2}}(\omega)$ for all $\omega\in\mathcal{M}$. Indeed the binary relation $\sim$ is an equivalence relation on $\mathcal{G}(\mathcal{M})$ (Zhou et al.,, 2025). If the geodesic $\gamma_{m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),m_{t}(\bm{Z}_{1,1:T_{0}})}$ serves as a substitute for the bias of the GSC estimator, the following result justifies the augmented GSC.
Proposition 3.1.
Under Assumption 3.2, suppose $\gamma_{m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),m_{t}(\bm{Z}_{1,1:T_{0}})}\sim\gamma_{Y_{1,t}^{(\mathrm{GSC})},Y_{1,t}(N)}$ for $t=T_{0}+1,\dots,T$. Then
@@
Figure 2 illustrates the augmented GSC. For Euclidean outcomes, the augmented GSC estimator can be represented as
$Y_{1,t}^{(\mathrm{AGSC})}$ $=\underbrace{\sum_{j=2}^{J+1}\overline{w}_{j}Y_{j,t}}_{=Y_{1,t}^{(\mathrm{GSC})}}+\overbrace{\widehat{m}_{t}(\bm{Z}_{1,1:T_{0}})-\underbrace{\sum_{j=2}^{J+1}\overline{w}_{j}\widehat{m}_{t}(\bm{Z}_{j,1:T_{0}})}_{=\widehat{m}_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})}}^{\text{bias correction}}=\widehat{m}_{t}(\bm{Z}_{1,1:T_{0}})+\sum_{j=2}^{J+1}\overline{w}_{j}\left(Y_{j,t}-\widehat{m}_{t}(\bm{Z}_{j,1:T_{0}})\right),$
which coincides with the augmented SC estimator of Ben-Michael et al., (2021). Hence, the proposed augmented GSC estimator $Y_{1,t}^{(\mathrm{AGSC})}$ is a natural extension of the conventional augmented synthetic control estimator.
Remark 3.4 (Regression estimator).
To implement the augmented GSC with $X_{j,t}\in\mathcal{M}$, one can use the geodesic optimal transport regression (Zhu and Müller,, 2023) to model $m_{t}(\bm{Z}_{j,1:T_{0}})$. For the special case where the covariates are Euclidean with $X_{j,t}\in\mathbb{R}^{q}$, employing global Fréchet regression (Petersen and Müller,, 2019) is an option to implement $m_{t}(\bm{Z}_{j,1:T_{0}})=m_{t}(\bm{X}_{j,1:T_{0}})$ in order to predict $Y_{j,t}(N)$ for $t=T_{0}+1,\dots,T$, as follows:
$\widehat{m}_{t}(\bm{x})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{J}\sum_{j=2}^{J+1}\{1+(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})^{\prime}\widehat{\Sigma}_{1:T_{0}}^{-1}(\bm{x}-\bar{\bm{X}}_{1:T_{0}})\}d^{2}(\nu,Y_{j,t}),$
@@
$\bar{\bm{X}}_{1:T_{0}}=\frac{1}{J}\sum_{j=2}^{J+1}\bm{X}_{j,1:T_{0}},\ \widehat{\Sigma}_{1:T_{0}}=\frac{1}{J}\sum_{j=2}^{J+1}(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})^{\prime}.$
Figure 2. Illustration of the augmented geodesic synthetic control method. The circles symbolize random objects in the geodesic metric space $\mathcal{M}$ and the arrows depict geodesics that connect the objects. The geodesic corresponding to the dashed red arrow from $Y_{1,t}^{(\mathrm{GSC})}$ to $Y_{1,t}(N)$ represents the bias of the geodesic synthetic control method in the setting considered in Section 3.3. The solid red arrow represents the geodesic connecting $m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})$ and $m_{t}(\bm{Z}_{1,1:T_{0}})$ that serves as a substitute to assess the bias of the geodesic synthetic control estimator. The dashed red arrow from $m_{t}(\bm{Z}_{1,1:T_{0}})$ to $Y_{1,t}(N)$ represents the bias correction for the geodesic synthetic control estimator corresponding to the augmented geodesic control estimator.
Remark 3.5 (Examples of geodesic transport maps).
@@
We provide here specific geodesic transport maps that satisfy Assumption 3.2 for Examples 2.12.5 in Section 2. For the space of networks in Example 2.1, the geodesic corresponds to the line segment connecting the start and end points. Then the geodesic transport map is given as $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$.
@@
For the space of compositional data in Example 2.2, the geodesic transport map can be interpreted as a rotation of the point $\omega$ along the geodesic determined by $\alpha$ and $\beta$. The tangent vector for the geodesic $\gamma_{\alpha,\beta}$ is given by $v_{\alpha,\beta}=\beta-(\alpha^{\prime}\beta)\alpha$, whose magnitude and direction encode the geodesic length and directionality needed to move from $\alpha$ to $\beta$ along the sphere. The geodesic transport map is then defined as $\Gamma_{\alpha,\beta}(\omega)=\mathrm{Exp}_{\omega}\left(\theta_{\alpha,\beta}v/\|v\|\right)$, where $\theta_{\alpha,\beta}$ is the angle between $\alpha$ and $\beta$, and $v=v_{\alpha,\beta}-(\omega^{\prime}v_{\alpha,\beta})\omega$ is the projection of $v_{\alpha,\beta}$ onto the tangent space at $\omega$.
@@
For the space of SPD matrices in Example 2.3, the geodesic transport map induced by the Frobenius metric is the line segment connecting two points. Then the geodesic transport map is $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$. The geodesic transport map induced by the Log-Euclidean metric is $\Gamma_{\alpha,\beta}(\omega)=\exp\{\log\omega+(\log\beta-\log\alpha)\}$ and for the power metric family it is $\Gamma_{\alpha,\beta}(\omega)=[\omega^{p}+(\beta^{p}-\alpha^{p})]^{1/p}$ where $p>0$ is a constant denoting the power, $A^{p}=U\Lambda^{p}U^{\prime}$ and $U\Lambda U^{\prime}$ is the usual spectral decomposition of $A\in\mathrm{Sym}^{+}_{m}$. Let $\mathcal{L}^{+}_{m}$ be the set of lower triangular matrices with positive diagonal entries. For any $S\in\mathrm{Sym}^{+}_{m}$, there exists a unique $L\in\mathcal{L}^{+}_{m}$ such that $S=LL^{\prime}$ and a diffeomorphism $\mathscr{L}$ between $\mathrm{Sym}^{+}_{m}$ and $\mathcal{L}^{+}_{m}$. For any $L\in\mathcal{L}^{+}_{m}$, let $\lfloor L\rfloor$ be the strictly lower triangular part and $\mathbb{D}(L)$ the diagonal part of $L$. For $\alpha,\beta,\omega\in\mathcal{L}^{+}_{m}$, the geodesic transport map induced by the Log-Cholesky metric is $\Gamma_{\alpha,\beta}(\omega)=\lfloor\omega\rfloor+(\lfloor\beta\rfloor-\lfloor\alpha\rfloor)+\exp\left\{\log(\mathbb{D}(\omega))+(\log(\mathbb{D}(\beta))-\log(\mathbb{D}(\alpha))\right\}$.
@@
For the Wasserstein space in Example 2.4, the geodesic transport map is $\Gamma_{\alpha,\beta}=F^{-1}_{\beta}\circ F_{\alpha}$, where $F_{\alpha}$ and $F^{-1}_{\beta}$ are the cumulative distribution function of the measure $\alpha$ and the quantile function of the measure $\beta$, respectively. The resulting endpoint of the geodesic $\gamma_{\omega,\zeta}$ is then $F^{-1}_{\zeta}=F^{-1}_{\beta}\circ F_{\alpha}\circ F^{-1}_{\omega}$ where $F^{-1}_{\omega}$ and $F^{-1}_{\zeta}$ denote the quantile functions of the measures $\omega$ and $\zeta$.
@@
For the space of functional data in Example 2.5, the geodesic transport map induced by the $L^{2}$ metric is the line segment connecting two points. Then the geodesic transport map is $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$.
4. Implementation and Simulations
4.1. Implementation details
Algorithm 1 outlines the implementation steps for the proposed geodesic synthetic control method. The optimization problem in Step 1 of Algorithm 1 requires customized approaches that depend on the specific metric and geometry of the underlying metric space. For networks (Example 2.1), symmetric positive-definite matrices (Example 2.3), distributions (Example 2.4), and functions (Example 2.5), the optimization reduces to a straightforward convex quadratic optimization problem that can be easily implemented (Stellato et al.,, 2020).
For the case of compositional data (Example 2.2), the optimization becomes more challenging, as it involves minimizing a bilevel objective where the weighted Fréchet mean $Y_{2:J+1,t}^{(\bm{w})}$ is itself an argmin problem. Due to the complexity of analytical differentiation when nonlinear Fréchet means are involved, we adopt a derivative-free constrained optimization strategy. Specifically, we use the Constrained Optimization BY Linear Approximations (COBYLA) algorithm (Powell,, 1994), as implemented in the R package nloptr (Johnson,, 2008). The weighted Fréchet mean involved in the objective function is obtained using the R package manifold (Dai et al.,, 2021).
Input: data $\{Y_{j,t}:j=1,\dots,J+1,t=1,\ldots,T\}$.
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})})$
where $Y_{2:J+1,t}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t})$ for $t=1,\ldots,T_{0}$;
2 $Y_{1,t}^{(\mathrm{GSC})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,Y_{j,t})\longleftarrow$ the synthetic control unit;
Algorithm 1 Geodesic Synthetic Control Method
4.2. Simulations for networks and symmetric positive-definite matrices
For the network outcomes, we use graph Laplacians residing in the space $\mathcal{L}$ equipped with the Frobenius metric (see Example 2.1). Networks are simulated via a weighted stochastic block model with 10 nodes divided into two equal-sized communities. Edges within and between communities are generated using Bernoulli distributions with probabilities specified by the edge probability matrix $\begin{pmatrix}0.75&0.1\\ 0.1&0.75\end{pmatrix}$. The weights of existing edges are generated as $\sin(0.1\pi t)+e^{-0.1t}((0.1j-0.5)^{2}-\sin(0.1\pi t))$. The graph Laplacian associated with each network is then considered to be the outcome.
For outcomes in the space of SPD matrices $\mathrm{Sym}_{m}^{+}$ equipped with the Log-Euclidean metric (see Example 2.3), we generate the underlying SPD matrices $\mu$ and $U$ from a Wishart distribution with 12 degrees of freedom and a scale matrix of $0.1\mathbf{I}_{10}$, where $\mathbf{I}_{10}$ denotes the $10\times 10$ identity matrix. The deterministic time trend and the unit-specific latent characteristic are constructed as $\mu_{t}=0.1t\mu$ and $U_{j}=e^{(0.1j-0.5)^{2}}U$, respectively. The parameter $\alpha_{t}$ governing the geodesic interpolation is set to $\alpha_{t}=\log(0.1(t+1))$.
Figure 3. Simulated counterfactual outcome for the treated unit had it not received the treatment $Y_{1,T}(N)$ (left) and the corresponding synthetic control outcome $Y_{1,T}^{(\mathrm{GSC})}$ (right) using the proposed method for (A) networks and (B) SPD matrices. For networks, nodes from the two communities are represented by red circles and blue squares, respectively. For SPD matrices, outcomes are visualized as heatmaps.
5. Real Data Example
5.1. Impact of the 2011 Great East Japan Earthquake on Employment-Population Ratio by Industry
We analyze the impact of the 2011 Great East Japan Earthquake on the employment-population ratio by employment category, using data from the Research Institute of Economy, Trade and Industry (RIETI) database (https://www.rieti.go.jp/jp/database/R-JIP2021/index.html). These data provide annual employment figures by industry for each prefecture in Japan from 1994 to 2018. Employment is categorized into three broad groups: primary (agriculture, forestry, and fisheries), secondary (manufacturing, electricity, gas, water supply, waste management and construction) and tertiary (services, including wholesale/retail, finance, and healthcare). For each prefecture and each year, the employment shares (proportions) across these three categories are represented as a three-dimensional compositional vector. The component-wise square root transformation (see Example 2.2) maps these compositional outcomes onto the positive quadrant of the sphere $\mathcal{S}_{+}^{2}\subset\mathbb{R}^{3}$, which we endow with the geodesic metric. Unlike standard synthetic control approaches that typically analyze each compositional component separately, the proposed approach considers the entire compositional vector jointly. This holistic perspective reflects the intrinsic geometry of compositional data, thereby capturing interdependencies of employment across the three categories.
Figure 4. Ternary plot depicting employment compositions, where employment composition in the treated unit, Miyagi, is represented by circles, while employment compositions in the control prefectures are shown as triangles. Different years are distinguished by color. The synthetic controls for the post-treatment periods 2011, 2012 and 2013 are shown as hollow circles and are distinguished from the observed employment compositions for these years by a similar proportion of primary category employment, an increased proportion of secondary category employment and a reduced proportion of tertiary category employment.
5.2. Impact of the 1972 Abortion Legislation on Fertility Patterns in East Germany
The study of fertility rates has attracted substantial attention, with an increasing emphasis on analyzing age-specific fertility rates (ASFR) rather than relying solely on summary measures such as total fertility rate (Chen et al.,, 2017). Represented as fertility curves across reproductive ages, ASFR provides richer insights into fertility behaviors, allowing researchers to detect age-specific causal effects that aggregate statistics alone cannot capture. For example, a policy intervention might disproportionately influence fertility behaviors among younger women, information that is lost when relying solely on aggregate metrics.
In March 1972, East Germany introduced legislation permitting women to terminate pregnancies within the first twelve weeks, significantly liberalizing abortion access. Although the primary intention was reproductive autonomy, an unintended consequence was a substantial drop in fertility, with the total fertility rate decreasing notably from 2.1 children per woman in 1971 to 1.5 by 1975 (Büttner and Lutz,, 1990). To gain deeper insights into how this policy affected fertility behavior across different reproductive ages, we applied the proposed GSC to ASFR curves. When viewed as functions over reproductive ages, these are functional data in the space $L^{2}$ (see Example 2.5).
The data for this analysis were obtained from the Human Fertility Database,, which provides ASFR annually for ages 12 to 55 for each calendar year across different countries. We select East Germany as the treated unit, given its distinct policy change, and utilize a control group composed of 26 countries that did not experience comparable abortion liberalization during this period (Austria, Belgium, Bulgaria, Belarus, Canada, Switzerland, Czechia, Denmark, Spain, Estonia, Finland, France, Hungary, Iceland, Ireland, Italy, Japan, Lithuania, Netherlands, Portugal, Russia, Slovakia, Sweden, Ukraine, United Kingdom (England and Wales), and United States). We define 1965–1971 as pre-treatment periods and 1972–1975 as post-treatment periods, so that it is possible to capture both immediate and medium-term policy impacts. Figure 5 visualizes the ASFR curves for East Germany and the control countries from 1966 to 1975 using heat maps. A clear reduction in fertility rates is evident for East Germany during the post-treatment period (1972–1975) but other countries also experience reductions in fertility between 1965 and 1975.
Figure 5. Age-specific fertility rates (ASFR) for East Germany and the control countries from 1966 to 1977 shown as heat maps. For each of the countries, each column represents the ASFR across ages 12 to 55 for a given calendar year that is indicated at the bottom of the figure.
Figure 6. Heat maps of age-specific fertility rates (ASFR) for East Germany (left) and the synthetic control unit (right) from 1966 to 1975. Each column represents the ASFR across ages 12 to 55 (vertical axis) for each calendar year (horizontal axis).
6. Extension to Geodesic Synthetic Difference-in-Differences
6.1. Construction of the estimator

### v2

@@ 7 Data Availability Statement
Let $(\mathcal{M},d)$ be a unique geodesic space. For any two points $\alpha,\beta\in\mathcal{M}$, there exists a geodesic transport map $\Gamma_{\alpha,\beta}:\mathcal{M}\to\mathcal{M}$ such that:
  1. (i)
    $\Gamma_{\alpha,\beta}(\alpha)=\beta$, i.e., the map sends the starting point $\alpha$ to the endpoint $\beta$;
  2. (ii)
    for any $\omega\in\mathcal{M}$, there exists a unique $\zeta\in\mathcal{M}$ such that $\Gamma_{\alpha,\beta}(\omega)=\zeta$, i.e., the displacement from $\alpha$ to $\beta$ can be consistently applied to any other point $\omega$.
This assumption ensures that the geodesic segment connecting $\alpha$ and $\beta$ can be “extended” from any other point $\omega\in\mathcal{M}$, producing a uniquely defined endpoint $\zeta$. In other words, the direction and length of the geodesic from $\alpha$ to $\beta$ induce a transformation that can be applied uniformly to all points in the space. In Euclidean space $\mathbb{R}^{q}$, this notion reduces to the familiar vector translation:
$\Gamma_{\alpha,\beta}(\omega)\;=\;\omega+(\beta-\alpha),$
so that each point $\omega$ is shifted by the same displacement vector that moves $\alpha$ to $\beta$. This construction extends naturally to general vector spaces, where geodesics correspond to linear displacements, and to Riemannian manifolds, where it is realized through parallel transport along geodesics (Yuan et al., 2012; Lin and Yao, 2019; Chen et al., 2023a).
The concept of geodesic transport maps is broad and applies to many of the spaces considered in this paper, including networks, compositional data, covariance matrices, distributions, and functions; see Section S.1.1 for more detail.
@@
Consider a metric space $(\mathcal{M}_{X},d_{X})$ that may differ from $(\mathcal{M},d)$ and the corresponding data pairs $\{(\bm{Y}_{j,1:T},\bm{X}_{j,1:T_{0}})\}_{1\leq j\leq J+1}$, where $\bm{Y}_{j,1:s}=(Y_{j,1},\dots,Y_{j,s})\in\mathcal{M}^{s}$ and $\bm{X}_{j,1:s}=(X_{j,1},\dots,X_{j,s})\in\mathcal{M}_{X}^{s}$. Denote by $m_{t}(\bm{Z}_{j,1:T_{0}})$ a regression model aiming to predict $Y_{j,t}(N)$ for $t=T_{0}+1,\dots,T$, where $\bm{Z}_{j,1:T_{0}}=(\bm{Y}_{j,1:T_{0}},\bm{X}_{j,1:T_{0}})$ and let $\hat{m}_{t}(\cdot)$ be its estimated version using data $\{(Y_{j,t},\bm{Z}_{j,1:T_{0}})\}_{2\leq j\leq J+1}$. Define
$m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}\overline{w}_{j}d^{2}(\nu,m_{t}(\bm{Z}_{j,1:T_{0}})),$
@@
$Y_{1,t}^{(\mathrm{AGSC})}$ $=\Gamma_{\hat{m}_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),\hat{m}_{t}(\bm{Z}_{1,1:T_{0}})}(Y_{1,t}^{(\mathrm{GSC})}).$
To see the rationale for this estimator, denote by $\mathcal{G}(\mathcal{M})=\{\gamma_{\alpha,\beta}:\alpha,\beta\in\mathcal{M}\}$ the space of geodesics of the unique geodesic space $\mathcal{M}$. To compare geodesics, we define a binary relation $\sim$ on $\mathcal{G}(\mathcal{M})$ through the geodesic transport map, where $\gamma_{\alpha_{1},\beta_{1}}\sim\gamma_{\alpha_{2},\beta_{2}}$ if and only if $\Gamma_{\alpha_{1},\beta_{1}}(\omega)=\Gamma_{\alpha_{2},\beta_{2}}(\omega)$ for all $\omega\in\mathcal{M}$. Indeed the binary relation $\sim$ is an equivalence relation on $\mathcal{G}(\mathcal{M})$ (Zhou et al., 2025). If the geodesic $\gamma_{m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),m_{t}(\bm{Z}_{1,1:T_{0}})}$ serves as a substitute for the bias of the GSC estimator, the following result justifies the augmented GSC.
Proposition S1.
Under Assumption S1, suppose $\gamma_{m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}}),m_{t}(\bm{Z}_{1,1:T_{0}})}\sim\gamma_{Y_{1,t}^{(\mathrm{GSC})},Y_{1,t}(N)}$ for $t=T_{0}+1,\dots,T$. Then
@@
Figure S1: Illustration of the augmented geodesic synthetic control method. The circles symbolize random objects in the geodesic metric space $\mathcal{M}$ and the arrows depict geodesics that connect the objects. The geodesic corresponding to the dashed red arrow from $Y_{1,t}^{(\mathrm{GSC})}$ to $Y_{1,t}(N)$ represents the bias of the geodesic synthetic control method in the setting considered in Section S.1. The solid red arrow represents the geodesic connecting $m_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})$ and $m_{t}(\bm{Z}_{1,1:T_{0}})$ that serves as a substitute to assess the bias of the geodesic synthetic control estimator. The dashed red arrow from $m_{t}(\bm{Z}_{1,1:T_{0}})$ to $Y_{1,t}(N)$ represents the bias correction for the geodesic synthetic control estimator corresponding to the augmented geodesic control estimator.
Figure S1 illustrates the augmented GSC. For Euclidean outcomes, the augmented GSC estimator can be represented as
$Y_{1,t}^{(\mathrm{AGSC})}$ $=\underbrace{\sum_{j=2}^{J+1}\overline{w}_{j}Y_{j,t}}_{=Y_{1,t}^{(\mathrm{GSC})}}+\overbrace{\hat{m}_{t}(\bm{Z}_{1,1:T_{0}})-\underbrace{\sum_{j=2}^{J+1}\overline{w}_{j}\hat{m}_{t}(\bm{Z}_{j,1:T_{0}})}_{=\hat{m}_{t}^{(\overline{\bm{w}})}(\bm{Z}_{2:J+1,1:T_{0}})}}^{\text{bias correction}}$
$=\hat{m}_{t}(\bm{Z}_{1,1:T_{0}})+\sum_{j=2}^{J+1}\overline{w}_{j}\left(Y_{j,t}-\hat{m}_{t}(\bm{Z}_{j,1:T_{0}})\right),$
which coincides with the augmented SC estimator of Ben-Michael et al. (2021). Hence, the proposed augmented GSC estimator $Y_{1,t}^{(\mathrm{AGSC})}$ is a natural extension of the conventional augmented synthetic control estimator.
Remark S1 (Regression estimator).
To implement the augmented GSC with $X_{j,t}\in\mathcal{M}$, one can use the geodesic optimal transport regression (Zhu and Müller, 2023) to model $m_{t}(\bm{Z}_{j,1:T_{0}})$. For the special case where the covariates are Euclidean with $X_{j,t}\in\mathbb{R}^{q}$, employing global Fréchet regression (Petersen and Müller, 2019) is an option to implement $m_{t}(\bm{Z}_{j,1:T_{0}})=m_{t}(\bm{X}_{j,1:T_{0}})$ in order to predict $Y_{j,t}(N)$ for $t=T_{0}+1,\dots,T$, as follows:
$\hat{m}_{t}(\bm{x})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{J}\sum_{j=2}^{J+1}\{1+(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})^{\prime}\hat{\Sigma}_{1:T_{0}}^{-1}(\bm{x}-\bar{\bm{X}}_{1:T_{0}})\}d^{2}(\nu,Y_{j,t}),$
@@
$\bar{\bm{X}}_{1:T_{0}}=\frac{1}{J}\sum_{j=2}^{J+1}\bm{X}_{j,1:T_{0}},\ \hat{\Sigma}_{1:T_{0}}=\frac{1}{J}\sum_{j=2}^{J+1}(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})(\bm{X}_{j,1:T_{0}}-\bar{\bm{X}}_{1:T_{0}})^{\prime}.$
S.1.1 Examples of geodesic transport maps
@@
We provide here specific geodesic transport maps that satisfy Assumption S1 for Examples 1–5 in Section 2. For the space of networks in Example 1, the geodesic corresponds to the line segment connecting the start and end points. Then the geodesic transport map is given as $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$.
@@
For the space of compositional data in Example 2, the geodesic transport map can be interpreted as a rotation of the point $\omega$ along the geodesic determined by $\alpha$ and $\beta$. The tangent vector for the geodesic $\gamma_{\alpha,\beta}$ is given by $v_{\alpha,\beta}=\beta-(\alpha^{\prime}\beta)\alpha$, whose magnitude and direction encode the geodesic length and directionality needed to move from $\alpha$ to $\beta$ along the sphere. The geodesic transport map is then defined as $\Gamma_{\alpha,\beta}(\omega)=\mathrm{Exp}_{\omega}\left(\theta_{\alpha,\beta}v/\|v\|\right)$, where $\theta_{\alpha,\beta}$ is the angle between $\alpha$ and $\beta$, and $v=v_{\alpha,\beta}-(\omega^{\prime}v_{\alpha,\beta})\omega$ is the projection of $v_{\alpha,\beta}$ onto the tangent space at $\omega$.
@@
For the space of SPD matrices in Example 3, the geodesic transport map induced by the Frobenius metric is the line segment connecting two points. Then the geodesic transport map is $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$. The geodesic transport map induced by the Log-Euclidean metric is $\Gamma_{\alpha,\beta}(\omega)=\exp\{\log\omega+(\log\beta-\log\alpha)\}$ and for the power metric family it is $\Gamma_{\alpha,\beta}(\omega)=[\omega^{p}+(\beta^{p}-\alpha^{p})]^{1/p}$ where $p>0$ is a constant denoting the power, $A^{p}=U\Lambda^{p}U^{\prime}$ and $U\Lambda U^{\prime}$ is the usual spectral decomposition of $A\in\mathrm{Sym}^{+}_{m}$. Let $\mathcal{L}^{+}_{m}$ be the set of lower triangular matrices with positive diagonal entries. For any $S\in\mathrm{Sym}^{+}_{m}$, there exists a unique $L\in\mathcal{L}^{+}_{m}$ such that $S=LL^{\prime}$ and a diffeomorphism $\mathscr{L}$ between $\mathrm{Sym}^{+}_{m}$ and $\mathcal{L}^{+}_{m}$. For any $L\in\mathcal{L}^{+}_{m}$, let $\lfloor L\rfloor$ be the strictly lower triangular part and $\mathbb{D}(L)$ the diagonal part of $L$. For $\alpha,\beta,\omega\in\mathcal{L}^{+}_{m}$, the geodesic transport map induced by the Log-Cholesky metric is $\Gamma_{\alpha,\beta}(\omega)=\lfloor\omega\rfloor+(\lfloor\beta\rfloor-\lfloor\alpha\rfloor)+\exp\left\{\log(\mathbb{D}(\omega))+(\log(\mathbb{D}(\beta))-\log(\mathbb{D}(\alpha))\right\}$.
@@
For the Wasserstein space in Example 4, the geodesic transport map is $\Gamma_{\alpha,\beta}=F^{-1}_{\beta}\circ F_{\alpha}$, where $F_{\alpha}$ and $F^{-1}_{\beta}$ are the cumulative distribution function of the measure $\alpha$ and the quantile function of the measure $\beta$, respectively. The resulting endpoint of the geodesic $\gamma_{\omega,\zeta}$ is then $F^{-1}_{\zeta}=F^{-1}_{\beta}\circ F_{\alpha}\circ F^{-1}_{\omega}$ where $F^{-1}_{\omega}$ and $F^{-1}_{\zeta}$ denote the quantile functions of the measures $\omega$ and $\zeta$.
@@
For the space of functional data in Example 5, the geodesic transport map induced by the $L^{2}$ metric is the line segment connecting two points. Then the geodesic transport map is $\Gamma_{\alpha,\beta}(\omega)=\omega+(\beta-\alpha)$.
S.2 Geodesic Synthetic Difference-in-Differences
S.2.1 Construction of the estimator